OpenStudy (bloodydiamond):

how to derive this equation r = a sin 4 (theta) from polar to rectangular?

OpenStudy (fibonaccichick666):

do you want to change it from polar to rectangular or take the derivative of the polar and change it to cartesian?

OpenStudy (fibonaccichick666):

also, is that r or y?

OpenStudy (bloodydiamond):

I just want it to change from polar to rectangular

OpenStudy (bloodydiamond):

r

OpenStudy (fibonaccichick666):

ohk, and well, first check out this site https://www.mathsisfun.com/polar-cartesian-coordinates.html

OpenStudy (bloodydiamond):

thanks

OpenStudy (fibonaccichick666):

and this, because the way your problem is given. You actually have a rose http://jwilson.coe.uga.edu/EMAT6680Fa09/Samples/Assignment%2011/Polar%20Equations.html

OpenStudy (tkhunny):

Is it \(\sin^{4}(\theta)\) or \(\sin(4\theta)\)?

OpenStudy (bloodydiamond):

sin(4 [\theta \])

OpenStudy (bloodydiamond):

lol

OpenStudy (freckles):

\[\sin(4 \theta)=2 \sin(2 \theta) \cos(2 \theta)\] use double angle identity for both \[\sin(2 \theta) \text{ and } \cos(2 \theta)\]

OpenStudy (freckles):

\[\text{ then use } \sin(\theta)=\frac{y}{r} \text{ and } \cos(\theta)=\frac{x}{r}\]

OpenStudy (bloodydiamond):

ok

OpenStudy (bloodydiamond):

tnx

OpenStudy (freckles):

np what did you get ?

OpenStudy (bloodydiamond):

still doing it

OpenStudy (freckles):

oh

OpenStudy (freckles):

the result is going to be pretty ugly looking

OpenStudy (freckles):

\[r = \sin ( 4 \theta) \\ r=2 \sin(2 \theta) \cos(2 \theta) \\ r= 2 (2 \sin(\theta) \cos(\theta)) (\cos^2(\theta)-\sin^2(\theta)) \\ r=4 \sin(\theta) \cos(\theta)(\cos^2(\theta)-\sin^2(\theta)) \\ r=4 \sin(\theta) \cos(\theta) ([\cos(\theta)]^2-[\sin(\theta)]^2)\] replace sin(theta) with y/r and replace cos(theta) with x/r

OpenStudy (freckles):

have you gotten this far?

OpenStudy (freckles):

if there really is a constant a in there we can throw in there afterwhile right now it is just going to be annoying

OpenStudy (bloodydiamond):

I did it the wrong way. i substitute the values first

OpenStudy (freckles):

how did you do that?

OpenStudy (bloodydiamond):

r = 2 (2y/r) (2x/r)

OpenStudy (bloodydiamond):

like that

OpenStudy (freckles):

\[r = \sin ( 4 \theta) \\ r=2 \sin(2 \theta) \cos(2 \theta) \\ r= 2 (2 \sin(\theta) \cos(\theta)) (\cos^2(\theta)-\sin^2(\theta)) \\ r=4 \sin(\theta) \cos(\theta)(\cos^2(\theta)-\sin^2(\theta)) \\ r=4 \sin(\theta) \cos(\theta) ([\cos(\theta)]^2-[\sin(\theta)]^2) \\ r=4 \frac{y}{r} \frac{x}{r}([\frac{x}{r}]^2-[\frac{y}{r}]^2)\]

OpenStudy (freckles):

\[r=\frac{4yx}{r^2}(\frac{x^2-y^2}{r^2}) \\ r=\frac{4yx(x^2-y^2)}{r^4}\]

OpenStudy (freckles):

\[\text{ recall } r^2=x^2+y^2 \]

OpenStudy (bloodydiamond):

ok

OpenStudy (freckles):

\[\text{ so } r^4=(x^2+y^2)^2 \\ \text{ and } r= \pm \sqrt{x^2+y^2}\]

OpenStudy (freckles):

\[\pm \sqrt{x^2+y^2}=\frac{4yx(x^2-y^2)}{(x^2+y^2)^2} \\ \text{ squared both sides so I didn't have to deal with the } \pm \\ x^2+y^2=(\frac{4y(x^2-y^2)}{(x^2+y^2)^2})^2\]

OpenStudy (freckles):

this is the rectangular equation for the polar equation r=sin(4 theta) just multiply that first one by in just the previous post to this one by a since you had r=a sin(4 theta) \[\pm \sqrt{x^2+y^2} =a \cdot \frac{ 4yx(x^2-y^2)}{(x^2+y^2)^2} \\ \text{ then \square both sides } \\ x^2+y^2=(a \frac{4 yx (x^2-y^2)}{(x^2+y^2)^2})^2\]

OpenStudy (freckles):

now if the question really was \[r=a \sin^4(\theta) \\ \text{ then multiply both sides by } r^4 \\ r^5 =a (r \sin(\theta))^4 \\ \text{ recall } \sin(\theta)=\frac{y}{r} \implies r \sin(\theta)=y \\ r^5=a(y)^4 \\ \text{ or } r^5=ay^4 \\ \text{ square both sides } (r^2)^5=a^2 y^8 \\ (x^2+y^2)^5=a^2 y^8\]

OpenStudy (freckles):

either way crazy equations there

OpenStudy (bloodydiamond):

wow thank you very much

OpenStudy (freckles):

np

OpenStudy (freckles):

were these for fun? or did your teacher really ask you to do these?

OpenStudy (freckles):

that last equation I did I don't know if wolfram has problems with the rectanglular form but here is the graph when I enter in polar form: http://www.wolframalpha.com/input/?i=r%3Dsin%5E4%28theta%29 here is the graph when I enter in rectangular form: http://www.wolframalpha.com/input/?i=%28x%5E2%2By%5E2%29%5E5%3Dy%5E8+to+polar+form notice they look very much similar but the second one seems a little jagged in some areas

OpenStudy (freckles):

anyways i have to go peace